Grade 9Mechanics

A ball is dropped on the floor from a height of 20m. It rebounds to a height of 1om. If the ball is in contact with floor for 0.1s. what is average acceleration during contact

Deepak kumar
7 Years agoGrade 9

2 Answers

jim larrington
7 Years ago
Use v^2=u^2+2as formula to find the velocity of the ball when it interacts with the floor. the answer is 26.3 metres per second.
Ashok Mishra
7 Years ago
u of the ball when it strikes the ground
=√2gh
=√2×10×20=20 m/s
 v of the ball when it bounces and reaches a max. height 10 m
=√2gh
=√2×10×10=10√2 m/s
avg acc^n = u-v/t
10√2-20/0.1 = -5.85 × 10^-3
ASKIITIANS EXPERT
SAMBHAV MISHRA